# Control variates for variance-reduced ratio of means estimators

### v1

@@ 6 Conclusion
In this paper, we introduced new optimized control variates coefficients for a variance-reduced ratio of means estimator. The resulting CV/CV estimator guarantees variance reduction when the coefficients $\alpha$ and $\beta$ are jointly optimized to minimize the overall variance. Significant variance reduction can be achieved, depending on the correlation structure between the variables of interest and the control variates. In particular, we showed that previously proposed coefficients could lead to a variance increase. To illustrate the approach, we applied the estimators with approximate control variates to estimate the strut mass fraction on a multi-fidelity aircraft design dataset.
Future work could aim to apply this method to reduce the variance of conditional mean estimators, such as the tail dependence [Malevergne and Sornette, 2002] or the conditional value at risk [Hong et al., 2014], as well as the self-normalized importance sampling estimator [Bugallo et al., 2017]. The method could also be generalized to estimators defined as functions of ratios. However, the current coefficients do not guarantee variance reduction in that setting; new coefficients would need to be derived for the full estimator.
@@
Acknowledgements
We gratefully acknowledge Eric Nguyen Van (ONERA/DTIS) for providing the dataset used in the application and for his valuable explanations of the aircraft design considerations behind it.
@@ Appendix A Deriving the optimal parameters and proof of variance reduction
To derive the optimal parameters and the variance reduction obtained with the estimators studied, an expression of the variance of a ratio is needed. Denote $X/Z$ the studied ratio, where $X$ and $Z$ are real random variables with non-zero means. Using first-order approximations and the delta method [Casella and Berger, 2024, example 5.5.27] [Sukhatme et al., 1970, section 4a.7] [Cochran, 1977, section 6.9], the variance of the Monte-Carlo estimator $\frac{\overline{X_{n}}}{\overline{Z_{n}}}$ of this ratio can be approximated by:
$\boxed{\text{Var}\left(\frac{\overline{X_{n}}}{\overline{Z_{n}}}\right)=\frac{1}{n}\left(\frac{1}{\mathbb{E}[Z]^{2}}\text{Var}(X)+\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[Z]^{4}}\text{Var}(Z)-2\frac{\mathbb{E}[X]}{\mathbb{E}[Z]^{3}}\text{Cov}\left(X,Z\right)\right)+\mathcal{O}(n^{-2})}$ (31)
@@ A.1 Variance of the MC/MC estimator
Let $(A_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $C$. The variance of the MC/MC estimator can be approximated as
@@ A.2.1 Variance of the CV/MC estimator
Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, and $C$, where $B$ is a control variate. The variance of the CV/MC estimator can be approximated as
@@ A.3.1 Variance of the CV/CV estimator
Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, $C$, and $D$, where $B$ and $D$ are control variates. The variance of the CV/CV estimator can be approximated as
@@
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $\approx\frac{1}{\mathbb{E}[\hat{C}_{CV,n}]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[\hat{A}_{CV}]^{2}}{\mathbb{E}[\hat{C}_{CV,n}]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[\hat{A}_{CV}]}{\mathbb{E}[\hat{C}_{CV,n}]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
@@ A.3.2 Computing the optimal coefficients of the CV/CV estimator
with $M=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})&-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ -2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\end{bmatrix}$ and $\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\\ 2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\end{bmatrix}$.
@@
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$ $\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\left(\eta+\beta_{o}R\right)^{2}\text{Var}(D)-2\left(\eta+\beta_{o}R\right)\text{Cov}(A,D)+R^{2}\text{Var}(C)+\beta_{o}^{2}R^{2}\text{Var}(D)\\ &-2\beta_{o}R^{2}\text{Cov}(C,D)-2R\text{Cov}(A,C)+2\left(\eta+\beta_{o}R\right)R\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)\\ &-2\left(\eta+\beta_{o}R\right)\beta_{o}R\text{Var}(D)\Big)\end{aligned}\right$.
@@ A.3.4 Comparing the variances of the MC/MC and the CV/CV estimators with different coefficients
$\bullet$ With the coefficients from [Gordon et al., 1982] $\alpha_{g}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ and $\beta_{g}=\frac{\text{Cov}(C,D)-\frac{1}{R}\text{Cov}(A,D)+\frac{\alpha}{R}\text{Cov}(B,D)}{\text{Var}(D)}$, the variance difference is:
@@ A.4 Calculs approximate CV
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT in order to improve language and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.

### v2

@@ 6 Conclusion
In this paper, we introduced new optimized control variates coefficients for a variance-reduced ratio of means estimator. The resulting CV/CV estimator guarantees variance reduction when the coefficients $\alpha$ and $\beta$ are jointly optimized to minimize the overall variance. Significant variance reduction can be achieved, depending on the correlation structure between the variables of interest and the control variates. In particular, we showed that previously proposed coefficients could lead to a variance increase. To illustrate the approach, we applied the estimators with approximate control variates to estimate the strut mass fraction on a multi-fidelity aircraft design dataset and conditional values-at-risk on a multi-fidelity electromagnetic dataset.
Future work could aim to apply this method to reduce the variance of conditional mean estimators, such as the tail dependence [malevergne_investigating_2002] or the self-normalized importance sampling estimator [bugallo_adaptive_2017]. The method could also be generalized to estimators defined as functions of ratios. However, the current coefficients do not guarantee variance reduction in that setting; new coefficients would need to be derived for the full estimator.
Acknowledgments
We gratefully acknowledge Eric Nguyen Van (ONERA/DTIS) for providing the aircraft design dataset and sharing his insights into the associated design considerations, and Isabelle Junqua (ONERA/DEMR) for providing the electromagnetic dataset along with her guidance on the relevant physical considerations.
@@ Acknowledgments
Data Availability Statement
The electromagnetic dataset and corresponding code to reproduce all figures are accessible on GitHub at https://github.com/LouB-N/Control-variates-for-variance-reduced-ratio-of-means-estimators.git. The aircraft design dataset is under consideration for publication; however, it can be provided to reviewers upon request.
Disclosure statement
The authors report there are no competing interests to declare.
Declaration of generative AI use in the manuscript preparation process
During the preparation of this work, the authors used Chat-GPT 5 in order to improve language and readability. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ Appendix A Deriving the optimal parameters and proof of variance reduction
To derive the optimal parameters and the variance reduction obtained with the estimators studied, an expression of the variance of a ratio is needed. Denote $X/Z$ the studied ratio, where $X$ and $Z$ are real random variables with non-zero means. Using first-order approximations and the delta method [casella_properties_2024, example 5.5.27] [sukhatme1970sampling, section 4a.7] [cochran_sampling_1977, section 6.9], the variance of the Monte-Carlo estimator $\frac{\overline{X_{n}}}{\overline{Z_{n}}}$ of this ratio can be approximated by:
$\boxed{\text{Var}\left(\frac{\overline{X_{n}}}{\overline{Z_{n}}}\right)=\frac{1}{n}\left(\frac{1}{\mathbb{E}[Z]^{2}}\text{Var}(X)+\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[Z]^{4}}\text{Var}(Z)-2\frac{\mathbb{E}[X]}{\mathbb{E}[Z]^{3}}\text{Cov}\left(X,Z\right)\right)+\mathcal{O}(n^{-\frac{3}{2}})}$ (31)
@@ A.1 Variance of the MC/MC estimator
Let $(A_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,C\in L^{2}$. The variance of the MC/MC estimator can be approximated as
@@ A.2.1 Variance of the CV/MC estimator
Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,B,C\in L^{2}$, where $B$ is a control variate. The variance of the CV/MC estimator can be approximated as
@@ A.3.1 Variance of the CV/CV estimator
Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,B,C,D\in L^{2}$, where $B$ and $D$ are control variates. The variance of the CV/CV estimator can be approximated as
@@
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$
$\approx\frac{1}{\mathbb{E}[\hat{C}_{CV,n}]^{2}}\text{Var}(\hat{A}_{CV})+\frac{\mathbb{E}[\hat{A}_{CV}]^{2}}{\mathbb{E}[\hat{C}_{CV,n}]^{4}}\text{Var}(\hat{C}_{CV,n})-2\frac{\mathbb{E}[\hat{A}_{CV}]}{\mathbb{E}[\hat{C}_{CV,n}]^{3}}\text{Cov}\left(\hat{A}_{CV},\hat{C}_{CV,n}\right)$
@@ A.3.2 Computing the optimal coefficients of the CV/CV estimator
with $M=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Var}(\overline{B_{n}})&-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})\\ -2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{D_{n}})&2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Var}(\overline{D_{n}})\end{bmatrix}$
and $\begin{bmatrix}c_{1}\\ c_{2}\end{bmatrix}=\begin{bmatrix}\frac{2}{\mathbb{E}[C]^{2}}\text{Cov}(\overline{A_{n}},\overline{B_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{B_{n}},\overline{C_{n}})\\ 2\frac{\mathbb{E}[A]^{2}}{\mathbb{E}[C]^{4}}\text{Cov}(\overline{C_{n}},\overline{D_{n}})-2\frac{\mathbb{E}[A]}{\mathbb{E}[C]^{3}}\text{Cov}(\overline{A_{n}},\overline{D_{n}})\end{bmatrix}$.
@@
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)$
$\approx\frac{1}{n\mathbb{E}[C]^{2}}\left.\begin{aligned} \Big(&\text{Var}(A)+\left(\eta+\beta_{o}R\right)^{2}\text{Var}(D)-2\left(\eta+\beta_{o}R\right)\text{Cov}(A,D)+R^{2}\text{Var}(C)+\beta_{o}^{2}R^{2}\text{Var}(D)\\ &-2\beta_{o}R^{2}\text{Cov}(C,D)-2R\text{Cov}(A,C)+2\left(\eta+\beta_{o}R\right)R\text{Cov}(C,D)+2\beta_{o}R\text{Cov}(A,D)\\ &-2\left(\eta+\beta_{o}R\right)\beta_{o}R\text{Var}(D)\Big)\end{aligned}\right$.
@@ A.3.4 Comparing the variances of the MC/MC and the CV/CV estimators with different coefficients
$\bullet$ With the coefficients from [gordon_efficient_1982] $\alpha_{g}=\frac{\text{Cov}(A,B)}{\text{Var}(B)}$ and $\beta_{g}=\frac{\text{Cov}(C,D)-\frac{1}{R}\text{Cov}(A,D)+\frac{\alpha}{R}\text{Cov}(B,D)}{\text{Var}(D)}$, the variance difference is: